OpenStudy (shelby1290):

The profit P(x) of a cosmetics company, in thousands of dollars, is given by P(x)=-5x^2+400x-2550, where X is the amount spent on advertising, in thousands of dollars a) Determine the maximum profit the company can make b) Determine the amount spent on advertising that will result in the maximum proift

OpenStudy (anonymous):

they are asking for the second and first coordinate of the vertex respectively

OpenStudy (anonymous):

the first coordinate of \[y=ax^2+bx+c\] is \[-\frac{b}{2a}\] and the second coordinate is what you get when you replace x by the first coordinate

OpenStudy (shelby1290):

@satellite73 to find the maximum profit, don't i have to use the quadratic formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait hold on a second

OpenStudy (anonymous):

there is a minus sign in front right?\[ P(x)=-5x^2+400x-2550\]

OpenStudy (anonymous):

other wise it would not have a maximum, it would have a minimum to set it equal to zero and solve would tell you when the profit was 0, not a maximum

OpenStudy (shelby1290):

@satellite73 yes it is a -5 but the question is asking to find the maximum profit

jimthompson5910 (jim_thompson5910):

From P(x) = -5x^2 + 400x - 2550, we see that a = -5 and b = 400 Plug these values into h = -b/(2a) to get h = -b/(2a) h = -400/(2*(-5)) h = ______ (fill in the blank)

OpenStudy (shelby1290):

@jim_thompson5910 h=40

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

so the x coordinate of the vertex is x = 40 plug x = 40 into the P(x) function to find the y coordinate of the vertex

OpenStudy (shelby1290):

P(x) = -5x^2 + 400x - 2550 P(40) = -5(40)^2 + 400(40) - 2550 = -8,000+16,000 -2550 = 5450 @jim_thompson5910 like this?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

x is in thousands, so is P(x) so x = 40 means 40,000 dollars is spent on ads yielding a profit of 5450*1000 = 5,450,000 dollars, which is the max profit